OpenStudy (anonymous):

Let f be a linear function where f(2)= -5 and f(-3)=1. Find f(x). HOW TO DO?!?!?!?!?!

OpenStudy (raden):

the general of a linear function is f(x) = ax + b (with a and b are the constant) now given f(2) = -5, it means a(-2) + b = -5 or -2a + b = -5 ........ (1) then known the value of f(-3) = 1, it means a(-3) + b = 1 or -3a + b = 1 ........ (2) now you have to solve for a and b respectively. (in the 1st and the 2nd eq) you can solve them by elimination and subtitution method

OpenStudy (raden):

got it ?

OpenStudy (anonymous):

I still do not understand sir.

OpenStudy (anonymous):

They gave u two points that are on the line, so you can calculate the slope

OpenStudy (anonymous):

do u know how?

OpenStudy (anonymous):

what are the points..?

OpenStudy (anonymous):

They told u that f(2)= -5, which means when you plug in 2 as x, you get -5 as y, so the point (2,-5) is on the line

OpenStudy (anonymous):

okay and the other point is (-3,1), right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

u know the formula for slope?

OpenStudy (anonymous):

You mean the point slope formula?

OpenStudy (anonymous):

You're gonna need the slope to use that formula

OpenStudy (anonymous):

have u ever heard of rise over run?

OpenStudy (anonymous):

ohh i know is it (y-y1)/( x-x1)?

OpenStudy (anonymous):

Ya thats right, so find the slope and then you can use the slope and point on the line to find the point slope form

OpenStudy (raden):

the slope formula doesnt works here. like i said above we have 2 equations : -2a + b = -5 -3a + b = 1 now, can you solve for a and b, first ?

OpenStudy (anonymous):

Wait, i thought i was suppose to find the slope and then point slope formula to find the point slope form. What does a and b have to do with this?

OpenStudy (raden):

they are the constant, and again like i said above : the general of a linear function is f(x) = ax + b so, we need the values of a and b, after you geot them then just plug into equation above